Circular No. 7218

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990705 G. Celidonio, G. Tarei, and S. Rebecchi, BeppoSAX Scientific Opera-

tion Center, Telespazio, Rome; J. Heise and J. in’t Zand, Space Research Organization of the Netherlands, Utrecht; N. Masetti, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; F. Frontera, Universitá di Ferrara; and P. Soffitta, M. Feroci, and G. Gandolfi, Istituto di Astrofisica Spaziale, CNR, Rome, write: “A γ-ray burst with a bright x-ray counterpart triggered the BeppoSAX Gamma-Ray Burst Monitor (GRBM) on July 5.66765 UT and was simultaneously detected in the Wide Field Camera (WFC) unit 2. The burst lasted for ∼ 45 s in the 2–26-keV WFC data and shows a complex, multipeaked structure. The WFC peak intensity corresponds to a flux of ∼ 4 Crab. The 40–700-keV GRBM flux is very strong and has a similar duration, with a peak of 4530 counts/s. The WFC-determined position of the x-ray counterpart to GRB 990705 is α = 5^{h}09^{m}52^{s}, δ = −72^{o}08′.0 (equinox 2000.0), with an error radius of 3′ (99-percent confidence level). A follow-up observation with BeppoSAX narrow-field instruments is still in progress.”

NOVA IN M31 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent nova in M31 (NGC 224). The object was discovered and confirmed on unfiltered CCD frames obtained on July 1.4 and 2.4 UT (mag ∼ 16.8 on both images). The new star is located at α = 0^{h}42^{m}49^{s}.5, δ = +41^{o}15′06′′ (equinox 2000.0), which is 60′′ east and 63′′ south of the nucleus of M31. A KAIT image of the same field on 1998 July 23.5 (limiting mag ∼ 19.0) showed nothing at the position of the nova. A finding chart

is located at http://astron.berkeley.edu/~bait/1999/n31 1.html.

COMET C/1998 T1 (LINEAR) Visual m_{1} estimates: June 13.53 UT, 9.4 (N. Biver, Oahu, HI, 0.26-m

reflector); 23.74, 8.6 (S. T. Rae, Palmerston North, New Zealand, 0.11-m refl.); July 2.96, 8.9 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m refl.); 4.42, 8.8 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 5.48, 9.2 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m refl.).

1999 July 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green